OpenStudy (anonymous):

If $200 is invested at 3.5% simple interest for 6 years, what will be the total amount of money in the investment account at the end of the 6 years?

jimthompson5910 (jim_thompson5910):

A=P+Prt A=200+200*0.035*6 A=200+200*0.21 A=200+42 A=242 So if you invest $200 at an interest rate of 3.5%, you'll have $242 in the account after 6 years.

OpenStudy (anonymous):

A member of a credit union was loaned $2600 to pay for car repairs. The loan was made for 4 years at a simple interest rate of 5.5%. Find the interest due on the loan.

jimthompson5910 (jim_thompson5910):

Interest = Principal * Rate * Time I = Prt I = 2600*0.055*4 I = 572 So the interest due is $572

OpenStudy (anonymous):

I=Prt = 2600*0.055*4=572

OpenStudy (anonymous):

total amount =3172

OpenStudy (anonymous):

If $490 is invested in an account that earns 8%, compounded annually, what will the account balance be after 5 years? (Round your answer to the nearest cent.)

jimthompson5910 (jim_thompson5910):

A=P(1+r/n)^(n*t) A=490(1+0.08/1)^(1*5) A=490(1+0.08)^(1*5) A=490(1.08)^(1*5) A=490(1.08)^(5) A=490(1.4693280768) A=719.970757632 A=719.97 So the account will balance after 5 years will be $719.97